### v1

# Mental-Perceiver: Audio-Textual Multimodal Learning for Mental Health Assessment

Abstract
Mental disorders, such as anxiety and depression, have become a global issue that affects the regular lives of people across different ages. Without proper detection and treatment, anxiety and depression can hinder the sufferer’s study, work, and daily life. Fortunately, recent advancements of digital and AI technologies provide new opportunities for better mental health care and many efforts have been made in developing automatic anxiety and depression assessment techniques. However, this field still lacks a publicly available large-scale dataset that can facilitate the development and evaluation of AI-based techniques. To address this limitation, we have constructed a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) on anxiety and depression assessment of Mandarin-speaking adolescents. The MMPsy contains audios and extracted transcripts of responses from automated anxiety or depression assessment interviews along with the self-reported anxiety or depression evaluations of the participants using standard mental health assessment questionnaires. Our dataset contains over 7,700 post-processed recordings of interviews for anxiety assessment and over 4,200 recordings for depression assessment. Using this dataset, we have developed a novel deep-learning based mental disorder estimation model, named Mental-Perceiver, to detect anxious/depressive mental states from recorded audio and transcript data. Extensive experiments on our MMPsy and the commonly-used DAIC-WOZ datasets have shown the effectiveness and superiority of our proposed Mental-Perceiver model in anxiety and depression detection. The MMPsy dataset will be made publicly available later to facilitate the research and development of AI-based techniques in the mental health care field.
Introduction
Anxiety and depression are two common mental health disorders that significantly impact patients, characterized by a cluster of similar signs and symptoms. Individuals suffering from anxiety or depression often experience a persistent state of distress, excessive fear and worry, lack of interest in everyday activities, and diminished energy levels. Beyond emotional disturbances, people with anxiety or depression also exhibit various physiological symptoms, such as weight loss, appetite changes, insomnia, and menopause in female patients (Cowen, Harrison, and Burns 2012). Due to these psychological and physiological impacts, anxiety and depression can hinder their sufferers’ study, work, and daily life. These mental disorders have become a thorny issue in our modern society. According to a 2019 report by the World Health Organization (WHO) (de la Santé 2019), over 301 million people, including 58 million children and adolescents, suffer from anxiety disorders, while more than 280 million people, including 23 million children and adolescents, live with depression. Despite the widespread prevalence, these disorders are frequently under- or mis-diagnosed and the treatment rates remain low (Kessler 2012). This is largely due to the time-consuming and costly nature of diagnosis and treatment. Additionally, patients, particularly children and adolescents, may conceal their true mental state out of fear of stigma during mental health assessments such as self-report screenings or clinical interviews, leading to misdiagnosis.
Fortunately, recent advancements of digital and AI technologies provide new opportunities for better mental health care and many efforts have been made in developing automatic anxiety and depression assessment techniques (Mao, Wu, and Chen 2023; Hecker et al. 2022). Given the pioneering explorations and achievements in this field, however, several significant limitations hinder current research efforts. Firstly, many existing methods rely heavily on the expertise of psychologists during interviews for data collection, which incurs substantial costs and makes it difficult to construct large-scale annotated datasets. As a result, the size of current datasets related to anxiety and depression detection, such as DAIC-WoZ (Gratch et al. 2014) and AViD-Corpus (Valstar et al. 2014), is limited, typically involving only around 100 participants. Although Shen et al. (Shen, Yang, and Lin 2022) made efforts to develop the EATD-Corpus, an emotional audio-textual depression detection dataset utilizing the Self-Rating Depression Scale (SDS), their dataset still includes only 162 volunteers from a University. Besides, the effectiveness of the SDS questionnaires completed by these volunteers was not validated, casting doubt on the reliability of the results. Consequently, constructing larger datasets at a lower cost remains a significant challenge for enhancing automatic anxiety and depression detection systems.
In this work, we address key challenges in the study of automatic anxiety and depression detection, aiming to extend research into more realistic settings. First, we introduce a novel large-scale Multi-Modal Psychological assessment corpus (MMPsy) focused on anxiety and depression in Mandarin-speaking adolescents. The MMPsy corpus includes audio recordings and corresponding transcripts of responses from both anxious/depressed and non-anxious/non-depressed adolescent volunteers, collected using a human-computer interaction system. Self-reporting anxiety and depression rating scales were also collected and verified using the same system, which can be used as gold standard in training and evaluating machine-learning based techniques. After rigorous data preprocessing and cleaning, the corpus comprises 7,758 interview data points for anxiety detection and 4,266 for depression detection. To the best of our knowledge, MMPsy is the first publicly available adolescent psychological assessment corpus that simultaneously detects anxiety and depression while providing both audio and text data in Chinese.
In addition, we propose a novel mental disorder estimation network, termed Mental-Perceiver, to automatically detect anxious or depressive mental states based on users’ audio inputs and corresponding transcripts. The Mental-Perceiver leverages attention mechanisms to map multimodal inputs and category semantic priors to a fixed-size multimodal feature, using a learnable, sharable embedding for effective multimodal fusion. It then applies a deep, fully attentional network to process the fused multimodal feature for in-depth multimodal feature extraction. Finally, the network decodes the extracted feature using a learnable query array to produce the final mental state estimation. Extensive experiments conducted on the MMPsy corpus and the public DAIC-WOZ dataset demonstrate the effectiveness and superiority of our proposed Mental-Perceiver model.
Our contributions can be summarized as follows:
  • We introduce MMPsy, a large-scale multimodal psychological assessment corpus containing 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection.
  • We develop Mental-Perceiver, a novel mental disorder estimation network that automatically detects anxious or depressive mental states from audio and transcript data, utilizing a fully attentional network architecture for effective multimodal fusion.
  • Extensive experimental results on MMPsy and the public DAIC-WOZ dataset validate the effectiveness and superiority of our proposed Mental-Perceiver model in automatic mental disorder detection.
Related Work
Automatic Anxiety/Depression Detection
Early studies of automatic anxiety/depression detection were concentrated on extracting effective features from the responses to questions that were highly correlated with anxiety/depression. Williamson et al. (Williamson et al. 2016) used related semantic context cues entailed in the voice, video-based facial action units, and transcribed text of individuals and built a Gaussian Staircase Model to detect depression automatically. Yang et al. (Yang et al. 2016) selected depression-related questions after analyzing interview transcripts manually and constructed a decision tree with the selected questions to predict the participants’ depression states. Similarly, Sun et al. (Sun et al. 2017) first manually selected questions related to certain topics such as recent feelings and sleep quality by conducting content analysis based on the text transcripts of clinical interviews. Then, they extracted text features from these selected questions as the input of Random Forest to detect depression tendencies. Gong et al. (Gong and Poellabauer 2017) performed topic modeling to split the interviews into topic-related segments. Then, they maintained the most discriminating features by a feature selection algorithm. Giannakakis et al. (Giannakakis et al. 2017) focused on non-voluntary and semi-voluntary facial cues to estimate the emotion representation more objectively. They selected the most robust features from the features including eye-related events, mouth activity, head motion parameters, and heart rate estimated through camera-based photoplethysmography. Finally, they deployed a ranking transformation to investigate the correlation of facial parameters with a participant’s perceived amount of stress/anxiety.
With the advancement of deep learning, extracting and integrating multi-modal features through deep learning models have become particularly promising for anxiety/depression detection. Ma et al. (Ma et al. 2016) encoded the depressive audio characteristics with the combination of CNN and LSTM to predict the presence of depression. Yang et al. (Yang et al. 2017) trained a deep Convolution Neural Network (CNN)-based depression detection model with specially-designed a set of audio and video descriptors. Tuka et al. (Al Hanai, Ghassemi, and Glass 2018) selected audio features and text features that were strongly related to depression severity by computing Pearson Coefficients and built a Long Short-Term Memory (LSTM) network to assess depression tendency. Haque et al. (Haque et al. 2018) proposed a causal CNN model to summarize acoustic, visual, and linguistic features into embeddings which were then used to predict depressive states. Shen et al. (Shen, Yang, and Lin 2022) proposed a depression detection approach utilizing speech characteristics and linguistic contents from participants’ interviews. Lin et al. (Lin et al. 2022) used the biological information of speech, combined with deep learning to build a rapid binary classification model of depression in the elderly. Agarwal (Agarwal, Jindal, and Singh 2023) developed machine-learning solutions to diagnose anxiety disorders from audio journals of patients.
Anxiety/Depression Detection Datasets
The public anxiety/depression datasets are quite scarce (Gratch et al. 2014; Valstar et al. 2014; Shen, Yang, and Lin 2022) due to ethical issues. To the best of our knowledge, there are only three publicly available datasets  (Gratch et al. 2014; Valstar et al. 2014; Shen, Yang, and Lin 2022) referring to depression detection while there are no publicly available multi-modal datasets including audio, facial, or text referring to anxiety detection. The DAIC-WoZ dataset (Gratch et al. 2014) contains recordings and transcripts of 142 American participants who were clinically interviewed by a computer agent automatically. AVid-Corpus (Valstar et al. 2014) contains audios and videos of German participants answering a set of queries or reciting fables. However, the text transcripts are not provided. EATD-Corpus (Shen, Yang, and Lin 2022) is released to facilitate the research in depression detection. It consists of audios and text transcripts extracted from the interviews of 162 student volunteers recruited from Tongji University. Although EATD-Corpus provides the text transcripts, its data scale is small.
MMPsy: A New Benchmark
Since the public anxiety/depression datasets are quite scarce, we construct and release a new Chinese multi-modal mental health dataset named MMPsy to facilitate the research in both anxiety detection and depression detection. Like EATD-Corpus (Shen, Yang, and Lin 2022), MMPsy consists of audios and text transcripts extracted from the interviews of more than 20 thousand primary and secondary school students in the Guangdong Province of China. All the volunteers have signed informed consent and tried their best to guarantee the authenticity of all the information provided. Each volunteer is required to answer 10 specially designed questions about anxiety and depression and complete a GAD-7 questionnaire (Mossman et al. 2017) or PHQ-9 questionnaire (Kroenke, Spitzer, and Williams 2001) for different disorders detection. The GAD-7 questionnaire consists of 7 items that measure the severity of anxiety. It is a commonly used questionnaire for psychologists to screen the anxiety severity of individuals in practice. A raw GAD-7 score can be summarized from the questionnaire. For Chinese people, an index GAD-7 score greater than or equal to 10 implies that the individual is in anxiety. Similarly, the PHQ-9 questionnaire consists of 9 items that measure the severity of depression. It is a commonly used questionnaire for psychologists to screen the depression severity of individuals. A raw PHQ-9 score can be summarized from the questionnaire. For Chinese people, an index PHQ-9 score greater than or equal to 10 implies that the individual is in depression. According to the criterion, there are 704 anxious volunteers and 7,032 non-anxious volunteers in the anxiety subset of MMPsy while there are 853 depressed volunteers and 3,394 non-depressed volunteers in the depression subset of MMPsy. In the final dataset, the overall duration of response audios about anxiety is about 145.9 hours and the overall duration of response audios about depression is about 69.5 hours.
Parts Subset $\times$ $\checkmark$ Avg Duration (Sec)
Anxiety Train 5625 563 68.05
Validation 704 70 65.97
Test 703 71 68.67
Depression Train 2715 682 59.05
Validation 340 85 55.27
Test 339 86 61.76
Table 1: Statistics for our MMPsy dataset. $\times$ means the label is non-anxious or non-depressed and $\checkmark$ denotes the data label is anxious or depressed.
The construction of MMPsy can be summarized as two steps: data collection and data preprocessing.
  • Data collection. We develop a web app to conduct the interview and to collect audio responses and the questionnaires. Our web app will ask the interviewees 10 questions and let them finish the GAD-7 questionnaires or PHQ-9 questionnaires. During the interview, the audio responses will be recorded and uploaded to the server automatically. The results of the questionnaires also will be uploaded. In this step, we collected 17,247 raw interviews about anxiety and 11,306 raw interviews about depression. To ensure the authenticity of the questionnaires, we added some extra but irrelevant questions in the questionnaires, which are used to check the response reliability.
  • Data preprocessing. Since the raw interviews are noised, we perform several preprocessing operations on the collected data. First, we filtered out the unauthentic data by checking the responses of those extra questions in the questionnaires. Second, mute audios and audios less than 1 second are removed. Besides, we also removed the silent segments identified with voice activity detection in each recording. Third, the background noises are eliminated using Spleeter (Hennequin et al. 2020). After that, Paraformer (Gao et al. 2022) is deployed to extract textual transcripts from audios. To ensure the correctness of extracted transcripts, we checked and corrected the transcript manually by listening to the audio and comparing the consistency between the semantics of auidos and the transcripts at the word level.
After data preprocessing, we obtained 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection. We randomly split these data into a training set, a validation set, and a test set in an 8:1:1 ratio. In this way, the anxiety detection part of MMPsy consists of 6,188 training volunteers, 744 validation volunteers, and 744 test volunteers while the depression detection part of MMPsy contains 3,397 training volunteers, 425 validation volunteers, and 425 test volunteers. The data statistics are listed in the Table 1. For each volunteer, we pack their audios and transcripts according to their sequential order and resegment the data in 60 seconds with a 10-second overlap like the way in the work (Wei et al. 2022). Thus, the data of a volunteer can generate multiple samples.
Mental-Perceiver
Figure 1: Illustration of Mental-Perceiver.
This section describes the architecture of our proposed Mental-Perceiver as shown in Figure 1. We first encode by applying an attention module that maps multimodal input $x\in\mathbb{R}^{M\times D_{x}}$ to features in a latent space $z\in\mathbb{R}^{2\times D_{z}}$ by interacting with category prior $p\in\mathbb{R}^{2\times D_{p}}$ which is obtained by computing the center point of different text representations from different categories. Then, we conduct deep feature extraction on latent features $z$ by applying a series of attention modules that take in and return features $z^{\prime}$ in this latent feature space. Finally, we decode by applying an attention module that maps latent arrays $z^{\prime}$ and the query array $q\in\mathbb{R}^{2\times D_{q}}$to the final feature representation $y\in\mathbb{R}^{2\times D_{y}}$. Based on the final feature representation $y$, we apply a Linear layer to map the feature $y$ into class-wise logit outputs $y^{C_{0}}\in\mathbb{R}^{1\times 2}$ and $y^{C_{1}}\in\mathbb{R}^{1\times 2}$. $M$ is the length of multimodal input while $D_{x}$, $D_{z}$, $D_{q}$, and $D_{y}$ denote the feature dimension. With class-wise logits $y^{C_{0}}$ and $y^{C_{1}}$, to predict the class $c^{\prime}$ of the input $x$, we compute the mean values of corresponding vector elements in the $y^{C_{0}}$ and $y^{C_{1}}$ to obtain final logits $y^{\prime}$ followed by a Softmax function.
Basic Attention module
Following the pioneering work PerceiverIO (Jaegle et al. 2021), all attention modules deployed in our Mental-Perceiver are implemented as Transformer-style attention (Vaswani et al. 2017). Each attention module applies a global query-key-value (QKV) attention operation followed by a multi-layer perceptron (MLP). The MLP is independently applied to each element of the index dimension. Both the encoder and decoder take in two input feature arrays. The first array is used as input to the attention module’s key and value networks, and another array is used as input to the module’s query network for interacting and fusing with the first array. The attention module’s output has the same index dimension (the same number of elements) as the query input. Therefore, the attention module can be modeled as follows:
$\text{attention}(Q,K,V)=MLP(Softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V)$ (1)
Category Semantic Prior
Our Mental-Perceiver extracts deep features based on the latent features $z$ which is produced by fusing the multimodal input $x$ and the latent embedding $p$ with an attention module. $p$ is used as a query and $x$ is used as the key and value. This means that the output $z$ is the result that $p$ extracts and fuse semantic information from $x$. So, the initialization of the latent embedding $p$ is crucial for learning more discriminative features in the following steps to identify a user’s mental state according to the user’s audio and transcript text. Semantic prior has been shown to help greatly learn more discriminative features (Teney, Abbasnejad, and Hengel 2019; Dai et al. 2023; Ding et al. 2023). Therefore, we build semantic priors for different categories, one for the normal category and another for the mental disorder category. Then, we use these two semantic priors to initialize the latent embedding $p\in\mathbb{R}^{2\times D}$ with a learnable MLP for mapping the hidden size of semantic priors to $D$. In Mental-Perceiver, we fixed the parameters of these two semantic priors and only optimized the parameters of the learnable MLP semantic priors. To obtain the semantic priors, we simply compute the center points for different categories.
Formally, given the text representation set $\textbf{E}_{t}\in\mathbb{R}^{N\times H}$ of a category on the training set, where $N$ is the number of data samples for the current category and $H$ is the hidden size of a text representation, we can compute the semantic prior $p^{C_{i}}\in\mathbb{R}^{1\times H}$ by averaging the normalized $\textbf{E}_{t}$ at the index dimension. This procedure can be modeled as follows:
$p^{C_{i}}=Avg(Norm(\textbf{E}_{t}))$ (2)
where $Avg$ is the averaging function and $Norm$ is the z-score normalization function that normalizes each vector separately in $\textbf{E}_{t}$. Let $C_{0}$ and $C_{1}$ denote the classes of normal people and psychological patients, respectively, we first obtain two semantic priors $p^{C_{0}}$ and $p^{C_{1}}$ according to Equation (2). Then, we obtain $p$ by concatenating $p^{C_{0}}$ and $p^{C_{1}}$ at the index dimension followed by a learnable MLP layer as follows:
$p=MLP(p^{C_{0}}\odot p^{C_{1}})$ (3)
where $\odot$ denotes the concatenation.
Encoder
The encoder, consisting of an attention module, takes charge of mapping the multimodal input $x$ into latent features $z$ by utilizing category semantic prior-enhanced latent embedding $p$ to query the multimodal input $x$. This way will fuse the multimodal input $x$ and category semantic prior-enhanced latent embedding $p$ and build fused prior-guided latent features $z$, which will be used as input to the next deep feature extraction. The encoder can be modeled as follows:
$z=attention(p,x,x)$ (4)
Deep Feature Extraction
Once we obtain the latent features $z$, we conduct deep feature extraction based on the input latents $z$ by applying a series of attention modules that take in and return latents $z_{i}$ in this latent space iteratively.
This module can be modeled as follows:
$z_{1}$ $=attention(z,z,z)$ (5)
$z_{2}$ $=attention(z_{1},z_{1},z_{1})$
$z_{k}$ $=attention(z_{k-1},z_{k-1},z_{k-1})$
where $k$ is a hyper-parameter that is simply set to 8.
Decoder
The goal of the decoder is to produce a final class-wise logit output of size $2\times 2$, given a latent representation of size $2\times D_{z}$. Let $z^{\prime}=z_{k}$, the decoder first applies an attention module to map the latent $z^{\prime}$ to output features $y$. Then, the decoder applies a linear layer to map $y$ into the final class-wise logit output $[y^{C_{0}},y^{C_{1}}]\in\mathbb{R}^{2\times 2}$, where $y^{C_{0}}\in\mathbb{R}^{1\times 2}$ and $y^{C_{1}}\in\mathbb{R}^{1\times 2}$ are two logit outputs for indicating whether the multimodal input $x$ matches with semantic prior $p^{C_{0}}$ and $p^{C_{1}}$. Finally, we compute the mean vector of these two output logit vectors $y^{C_{0}}$ and $y^{C_{1}}$ at the index dimension as the final classification logits $y^{\prime}$. Therefore, the decoder can be modeled as follows:
$y$ $=attention(q,z^{\prime},z^{\prime})$ (6)
$[y^{C_{0}},y^{C_{1}}]$ $=Linear(y)$
$y^{\prime}$ $=Mean(y^{C_{0}}\odot y^{C_{1}})$
Training Objectives
To optimize our Mental-Perceiver, we deploy two losses. The first one is the matching loss $\mathcal{L}_{match}$ and another one is the classification loss $\mathcal{L}_{cls}$. Both these two losses are binary cross-entropy loss functions. The matching loss aims to optimize the matching degree between the multimodal input $x$ and its corresponding category semantic prior while the classification loss optimizes the model to be able to identify the inherent mental state according to multimodal input $x$ with the help of category prior $p$ and category query $q$.
Formally, given a multimodal input $x$ and its class $C_{x}$ which can be 0 or 1, the training objectives can be modeled as follows:
$\mathcal{L}_{match}=$ $-C_{x}(logy^{C_{0}}_{0}+logy^{C_{1}}_{1})$ (7)
$-(1-C_{x})(logy^{C_{0}}_{1}+logy^{C_{1}}_{0})$
$\mathcal{L}_{cls}=$ $-C_{x}logy^{\prime}_{0}-(1-C_{x})logy^{\prime}_{1}$
$\mathcal{L}=$ $\mathcal{L}_{match}+\mathcal{L}_{cls}$
where $y^{C_{0}}_{0}$ and $y^{C_{0}}_{1}$ are the 0-th element and 1-th element in the probability distribution obtained from class-wise logits $y^{C_{0}}\in\mathbb{R}^{1\times 2}$ by Softmax function while $y^{C_{1}}_{0}$ and $y^{C_{1}}_{1}$ are the 0-th element and 1-th element in the probability distribution obtained from class-wise logits $y^{C_{1}}\in\mathbb{R}^{1\times 2}$ by Softmax function. Similarly, $y^{\prime}_{0}$ and $y^{\prime}_{1}$ are the two probabilities on normal class (0) and mental disorder class (1) that can be obtained by applying the Softmax function on $y^{\prime}$.
Experiments
Datasets
We conduct experiments on both two subsets of MMPsy for anxiety detection and depression detection. We use MMPsy-Anxiety and MMPsy-Depression to represent these two subsets. Besides, we also verify our Mental-Perceiver on the Distress Analysis Interview Corpus - Wizard of Oz (DAIC-WOZ) dataset (Gratch et al. 2014). DAIC-WOZ contains clinical interviews of 189 participants designed to support the diagnosis of psychological distress conditions such as anxiety, depression, and posttraumatic stress disorder (PTSD). During each interview, several data in different formats as well as modalities are recorded simultaneously. However, only the acoustic recordings and transcriptions are chosen in this work for a fair comparison. Moreover, the given GT is an eight-item Patient Health Questionnaire depression scale (PHQ-8), which indicates the severity of depression. A PHQ-8 Score $\geq$ 10 implies that the participant is undergoing a mental disorder.
Baselines
The main baselines to be compared are listed as follows:
  • SVM (Pedregosa et al. 2011): a robust shallow model capable of performing binary classification efficiently by using a kernel trick, representing the data only through a set of pairwise similarity comparisons between the original data points using a kernel function, which transforms them into coordinates in the higher dimensional feature space.
  • RandomForest (Pedregosa et al. 2011): it is a robust ensemble learning method for classification by constructing a multitude of decision trees at training time. For classification tasks, the output of the random forest is the class selected by most trees.
  • XGBoost (Chen and Guestrin 2016): it is a robust toolbox for classification via an optimized distributed gradient boosting.
  • NUSD (Wang, Ravi, and Alwan 2023): it is a deep model ECAPA-TDNN enhanced with a speaker disentanglement method that utilizes a non-uniform mechanism of adversarial SID loss maximization.
  • ConvLSTM (Wei et al. 2022): it is a Convolutional Bidirectional LSTM with a sub-attention mechanism for linking heterogeneous information.
  • PerceiverIO (Jaegle et al. 2021): it is a general-purpose architecture that handles multimodal data with fully attention design and a flexible querying mechanism.
For the shallow models SVM, RandomForest, and XGBoost, we provide the following audio features as input: F0 statistics (mean), log-energy, zero-crossing-rate, loudness, pitch period entropy, jitters, shimmers, harmonics-to-noise ratio, detrended fluctuation analysis, linear spectral coefficients-0, linear spectral frequencies-0, formants (F1), and amplitude Shannon entropy. All these features can be extracted by applying Surfboard  [^1] https://github.com/novoic/surfboard (Lenain et al. 2020). Besides, we also extract topic words by TF-IDF as text features for these shallow models. For the deep models, we use BERT (Devlin et al. 2019) to extract features and use Mel-spectrum to represent audio.
Datasets Methods Acc UAR Sens Spec Precision Recall F1 Precision Recall F1
Normal Anxious
MMPsy-Anxiety SVM 0.83 0.64 0.41 0.87 0.94 0.87 0.9 0.24 0.41 0.3
RandomForest 0.82 0.63 0.39 0.87 0.93 0.87 0.9 0.23 0.39 0.29
XGBoost 0.88 0.6 0.25 0.94 0.93 0.94 0.93 0.3 0.25 0.27
NUSD 0.79 0.51 0.17 0.85 0.91 0.85 0.88 0.1 0.17 0.13
ConvLSTM 0.83 0.55 0.21 0.9 0.92 0.9 0.91 0.18 0.21 0.19
PerceiverIO 0.83 0.74 0.63 0.84 0.95 0.84 0.90 0.29 0.63 0.40
Mental-Perceiver (Ours) 0.85 0.76 0.65 0.87 0.96 0.87 0.92 0.34 0.65 0.45
Normal Depressed
MMPsy-Depression SVM 0.76 0.71 0.63 0.79 0.89 0.79 0.84 0.43 0.63 0.51
RandomForest 0.76 0.63 0.41 0.85 0.85 0.85 0.85 0.41 0.41 0.41
XGBoost 0.78 0.64 0.41 0.88 0.85 0.88 0.87 0.46 0.41 0.43
NUSD 0.67 0.51 0.26 0.78 0.8 0.78 0.79 0.22 0.26 0.24
ConvLSTM 0.77 0.53 0.14 0.93 0.81 0.93 0.86 0.32 0.14 0.2
PerceiverIO 0.81 0.77 0.66 0.85 0.91 0.85 0.88 0.53 0.66 0.59
Mental-Perceiver (Ours) 0.85 0.79 0.69 0.89 0.92 0.89 0.9 0.61 0.69 0.64
Normal Depressed
DAIC-WOZ SVM 0.38 0.35 0.29 0.42 0.58 0.42 0.49 0.17 0.29 0.22
RandomForest 0.7 0.54 0.14 0.94 0.72 0.94 0.82 0.5 0.14 0.22
XGBoost 0.6 0.53 0.36 0.7 0.72 0.7 0.71 0.33 0.36 0.34
NUSD 0.55 0.46 0.21 0.7 0.68 0.7 0.69 0.23 0.21 0.22
ConvLSTM 0.4 0.57 0.88 0.22 0.83 0.22 0.35 0.3 0.88 0.45
PerceiverIO 0.7 0.58 0.29 0.88 0.74 0.88 0.81 0.5 0.29 0.36
Mental-Perceiver (Ours) 0.79 0.66 0.36 0.97 0.78 0.97 0.86 0.83 0.36 0.5
Table 2: Performance Comparison with our Mental-Perceiver and various baselines on MMPsy and DAIC-WOZ. The best result is highlighted in bold.
Metrics
In classification tasks, True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) from the confusion matrix are metrics used to measure the accuracy of model predictions. TP refers to instances where the model correctly predicted them as the positive class, while TN refers to instances where the model correctly predicted them as the negative class. On the contrary, FP represents cases where the model incorrectly labeled instances as positive, and FN denotes instances of the positive class that were incorrectly classified as negative.
Based on these concepts, several common performance metrics can be defined to assess the model’s performance:
  • Accuracy: Accuracy (Acc) represents the proportion of all predictions that are correctly predicted.
  • Recall: Recall measures the model’s ability to identify all true positive instances, that is, the proportion of actual positives that are correctly identified.
  • Precision: Precision denotes the proportion of samples predicted as positive by the model that are actually positive, focusing on the accuracy of positive predictions.
  • F1-Score: F1 Score ranges from 1 to 0, with a value closer to 1 indicating better model performance, particularly in scenarios where the distribution of positive and negative samples is imbalanced. The F1 Score serves as a more comprehensive evaluation metric under such conditions.
  • Sensitivity: Sensitivity (Sens), also termed true positive rate, is the ratio of positive predictions to the number of actual positives. It is hence identical to the recall of the positive class, Recall(1).
  • Specificity: Specificity (Spec) is the ratio of negative predictions to the number of actual negatives, and therefore identical to Recall(0).
  • UAR: Although Acc is often used to evaluate model performance, it suffers from the data imbalance issue. The stronger the imbalance the more accuracy tends to reflect the performance of the majority class since it is a weighted accuracy. For this reason in areas where relatively small data sets predominate, such as the bio-medical field or in the paralinguistics field in general, researchers usually prefer and report the Unweighted Average Recall (UAR), which is defined as the average of Sensitivity and Specificity.
    $UAR=\frac{Sensitivity+Specificity}{2}$ (8)
During these metrics, we use Accuracy (Acc), Unweighted Average Recall (UAR), Sensitivity (Sens), and Specificity (Spec) as the main evaluation metrics for evaluating overall model performance. Meanwhile, we also report the Recall, Precision, and F1-score separately for different categories.
Datasets Modalities Acc UAR Sens Spec Precision Recall F1 Precision Recall F1
Normal Anxious
MMPsy-Anxiety Audio 0.87 0.52 0.10 0.95 0.91 0.95 0.93 0.16 0.10 0.12
Text 0.85 0.74 0.61 0.88 0.96 0.88 0.92 0.34 0.61 0.43
Text+Audio 0.85 0.76 0.65 0.87 0.96 0.87 0.92 0.34 0.65 0.45
Normal Depressed
MMPsy-Depression Audio 0.76 0.61 0.35 0.86 0.84 0.86 0.85 0.39 0.35 0.37
Text 0.81 0.74 0.62 0.86 0.90 0.86 0.88 0.53 0.62 0.57
Text+Audio 0.85 0.79 0.69 0.89 0.92 0.89 0.9 0.61 0.69 0.64
Table 3: Ablation study on different modalities of our Mental-Perceiver on MMPsy. The best result is highlighted in bold.
Implementation Details
We use Pytorch [^2] http://pytorch.org to implement our framework on Linux with two NVIDIA RTX 4090 GPU cards. The feature dimension $D_{x}$ is set to 768 and other dimensions $D_{z}$, $D_{q}$, and $D_{y}$ are all set to 512. In each epoch, all training data is shuffled randomly and then cut into mini-batches. The text feature and audio feature are concatenated as the multimodal input. We deploy the AdamW (Loshchilov and Hutter 2017) optimizer for model optimization. We trained models for 200 epochs with an initial learning rate of 0.00003 and used LambdaLR to adjust the learning rate during training. The early stopping with patience 15 is deployed to accelerate training. We use the validation set for model selection and report the performance on the test set.
Experiment Results
Main Results
The experiment results of our Mental-Perceiver and various baselines on three datasets MMPsy-Anxiety, MMPsy-Depression, and DAIC-WOZ are shown in Table 2. From the results on different datasets, we can conclude as follows.
On the MMPsy-Anxiety dataset, our Mental-Perceiver outperforms baselines on Acc, UAR, and Sens while achieving competitive performance on Spec. This shows that our Mental-Perceiver can achieve better overall anxiety detection performance with a better trade-off between Sensitivity and Specificity. For different categories, we can observe that our Mental-Perceiver achieves relatively high performance in the normal category and achieves the best performance in the anxious category.
On MMPsy-Depression dataset, our Mental-Perceiver outperforms baselines on Acc, UAR, and Sens while achieving competitive performance on Spec. This shows that our Mental-Perceiver can achieve better overall depression detection performance with a better trade-off between Sensitivity and Specificity. For different categories, we can observe that our Mental-Perceiver achieves the best precision and F1-score in the normal category while outperforming all baselines in the depressed category on all three metrics.
On DAIC-WOZ dataset, a similar conclusion can be reached. The Mental-Perceiver outperforms baselines on Acc, UAR, and Spec. Although the Mental-Perceiver’s sensitivity is lower than the baseline ConvLSTM, ConvLSTM is very poor in specificity, indicating that there is a high rate of misdiagnosis on ConvLSTM. According to the UAR, we can observe Mental-Perceiver has a better balance between the false positive rate (1 - Specificity) and the false negative rate (1 - Sensitivity). Besides, according to the metrics on different categories, we can observe that the Mental-Perceiver achieves the best overall performance in both two categories.
Overall, our Mental-Perceiver can achieve the best overall performance on different datasets for different mental disorder detection, showing the effectiveness and universality of our Mental-Perceiver for detecting different mental disorders. Besides, different models can achieve varying degrees of performance, indicating the effectiveness and usability of our MMPsy dataset as a benchmark for developing and evaluating mental disorder detection models.
Ablation study on different modalities
To verify the superiority of multimodal text-audio for mental disorder detection, we conduct an ablation study by using only audio, only text, and text+aduio as input to the Mental-Perceiver. The experimental results are shown in Table 3. We can observe that the Mental-Perceiver with multimodal inputs can achieve the best performance across various metrics on the MMPsy-Depression while achieving the best performance on UAR, Sensitivity, Precision in the normal category, Precision in the anxious category, Recall in the anxious category, and F1 on the anxious category and reasonable and competitive performance on Acc, Specificity, Recall on the normal category, and F1 on the normal category. Overall, multimodal input helps detect mental disorders.
Datasets Modalities Acc UAR Sens Spec
MMPsy-Anxiety Mental-Perceiver 0.85 0.76 0.65 0.87
-Category Prior 0.84 0.72 0.58 0.86
-Matching Loss 0.82 0.74 0.62 0.86
MMPsy-Depression Mental-Perceiver 0.85 0.79 0.69 0.89
-Category Prior 0.82 0.68 0.49 0.88
-Matching Loss 0.8 0.72 0.57 0.86
Table 4: Ablation study on the effects of category priors and match loss on MMPsy. The best result is highlighted in bold.
The effects of category priors and matching loss
To investigate the effects of the category priors and matching loss, we conduct a study by removing the category priors and matching loss. The results are shown in Table 4. It can be seen that each component can improve the performance across various metrics, showing the effectiveness of category priors and matching loss.
Conclusion
In this work, we construct a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in adolescents who use Mandarin. The MMPsy contains audios and extracted transcripts of responses from anxious/depressed and non-anxious/non-depressed adolescent volunteers. There are 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection. We further propose a novel mental disorder estimation network, named Mental-Perceiver, to detect anxious/depressive mental states automatically according to users’ audio and corresponding transcripts. Extensive experiments on MMPsy and the public DAIC-WOZ show the effectiveness and superiority of our proposed Mental-Perceiver.

### v2

@@
-Mental-Perceiver: Audio-Textual Multimodal Learning for Mental Health Assessment
+Mental-Perceiver: Audio-Textual Multi-Modal Learning for Estimating Mental Disorders
 Abstract
-Mental disorders, such as anxiety and depression, have become a global issue that affects the regular lives of people across different ages.
-Without proper detection and treatment, anxiety and depression can hinder the sufferer’s study, work, and daily life.
-Fortunately, recent advancements of digital and AI technologies provide new opportunities for better mental health care and many efforts have been made in developing automatic anxiety and depression assessment techniques.
-However, this field still lacks a publicly available large-scale dataset that can facilitate the development and evaluation of AI-based techniques.
-To address this limitation, we have constructed a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) on anxiety and depression assessment of Mandarin-speaking adolescents.
-The MMPsy contains audios and extracted transcripts of responses from automated anxiety or depression assessment interviews along with the self-reported anxiety or depression evaluations of the participants using standard mental health assessment questionnaires.
-Our dataset contains over 7,700 post-processed recordings of interviews for anxiety assessment and over 4,200 recordings for depression assessment.
-Using this dataset, we have developed a novel deep-learning based mental disorder estimation model, named Mental-Perceiver, to detect anxious/depressive mental states from recorded audio and transcript data.
-Extensive experiments on our MMPsy and the commonly-used DAIC-WOZ datasets have shown the effectiveness and superiority of our proposed Mental-Perceiver model in anxiety and depression detection.
-The MMPsy dataset will be made publicly available later to facilitate the research and development of AI-based techniques in the mental health care field.
+Mental disorders, such as anxiety and depression, have become a global concern that affects people of all ages.
+Early detection and treatment are crucial to mitigate the negative effects these disorders can have on daily life.
+Although AI-based detection methods show promise, progress is hindered by the lack of publicly available large-scale datasets.
+To address this, we introduce the Multi-Modal Psychological assessment corpus (MMPsy), a large-scale dataset containing audio recordings and transcripts from Mandarin-speaking adolescents undergoing automated anxiety/depression assessment interviews.
+MMPsy also includes self-reported anxiety/depression evaluations using standardized psychological questionnaires.
+Leveraging this dataset, we propose Mental-Perceiver, a deep learning model for estimating mental disorders from audio and textual data.
+Extensive experiments on MMPsy and the DAIC-WOZ dataset demonstrate the effectiveness of Mental-Perceiver in anxiety and depression detection.
+Datasetshttps://github.com/shuyeit/mmpsy-data
 Introduction
-Anxiety and depression are two common mental health disorders that significantly impact patients, characterized by a cluster of similar signs and symptoms.
-Individuals suffering from anxiety or depression often experience a persistent state of distress, excessive fear and worry, lack of interest in everyday activities, and diminished energy levels.
-Beyond emotional disturbances, people with anxiety or depression also exhibit various physiological symptoms, such as weight loss, appetite changes, insomnia, and menopause in female patients (Cowen, Harrison, and Burns 2012).
-Due to these psychological and physiological impacts, anxiety and depression can hinder their sufferers’ study, work, and daily life.
-These mental disorders have become a thorny issue in our modern society.
-According to a 2019 report by the World Health Organization (WHO) (de la Santé 2019), over 301 million people, including 58 million children and adolescents, suffer from anxiety disorders, while more than 280 million people, including 23 million children and adolescents, live with depression.
-Despite the widespread prevalence, these disorders are frequently under- or mis-diagnosed and the treatment rates remain low (Kessler 2012).
-This is largely due to the time-consuming and costly nature of diagnosis and treatment.
-Additionally, patients, particularly children and adolescents, may conceal their true mental state out of fear of stigma during mental health assessments such as self-report screenings or clinical interviews, leading to misdiagnosis.
-Fortunately, recent advancements of digital and AI technologies provide new opportunities for better mental health care and many efforts have been made in developing automatic anxiety and depression assessment techniques (Mao, Wu, and Chen 2023; Hecker et al.
-2022).
-Given the pioneering explorations and achievements in this field, however, several significant limitations hinder current research efforts.
-Firstly, many existing methods rely heavily on the expertise of psychologists during interviews for data collection, which incurs substantial costs and makes it difficult to construct large-scale annotated datasets.
-As a result, the size of current datasets related to anxiety and depression detection, such as DAIC-WoZ (Gratch et al.
+Anxiety and depression are prevalent mental disorders that can have a significant and detrimental impact on an individual’s life.
+If left untreated, these disorders can disrupt work, study, and social interactions.
+Symptoms include persistent negative emotions, behavioral changes, and physiological disturbances (Cowen, Harrison, and Burns 2012).
+However, misdiagnosis is common due to the overlap of symptoms with other physical and physiological conditions.
+The World Health Organization (WHO) 2019 report (de la Santé 2019) indicates that more than 301 million people, including 58 million children and adolescents, suffer from anxiety disorders.
+In addition, more than 280 million people, including 23 million children and adolescents, are affected by depression.
+Despite the prevalence of these conditions, the treatment rate for both anxiety and depression remains notably low due to the substantial time and financial commitments required for diagnosis and treatment (Kessler 2012).
+Furthermore, individuals, particularly children and adolescents, can conceal their true mental state during assessments and interviews, preventing an accurate diagnosis.
+To address this challenging issue, automatic detection of anxiety and depression offers a potential solution.
+The development of a suitable psychological assessment corpus related to anxiety and depression and the construction of an automated anxiety and depression detection system are essential.
+This will allow individuals to assess their anxious or depressive states privately and increase their willingness to consult with psychologists after conducting self-assessments.
+Furthermore, such a psychological assessment corpus and automated detection system would greatly assist psychologists in diagnosing anxiety and depression disorders, even when patients conceal their true mental states.
+A significant challenge facing current research in automatic anxiety and depression detection is the absence of extensive datasets to train and evaluate machine learning models.
+Existing methodologies frequently depend on the judgment of expert psychologists (e.g., during clinical interviews) for data collection, making the creation of large-scale annotated datasets costly.
+Publicly accessible datasets, such as DAIC-WOZ (Gratch et al.
 2014) and AViD-Corpus (Valstar et al.
-2014), is limited, typically involving only around 100 participants.
-Although Shen et al.
-(Shen, Yang, and Lin 2022) made efforts to develop the EATD-Corpus, an emotional audio-textual depression detection dataset utilizing the Self-Rating Depression Scale (SDS), their dataset still includes only 162 volunteers from a University.
-Besides, the effectiveness of the SDS questionnaires completed by these volunteers was not validated, casting doubt on the reliability of the results.
-Consequently, constructing larger datasets at a lower cost remains a significant challenge for enhancing automatic anxiety and depression detection systems.
-In this work, we address key challenges in the study of automatic anxiety and depression detection, aiming to extend research into more realistic settings.
-First, we introduce a novel large-scale Multi-Modal Psychological assessment corpus (MMPsy) focused on anxiety and depression in Mandarin-speaking adolescents.
-The MMPsy corpus includes audio recordings and corresponding transcripts of responses from both anxious/depressed and non-anxious/non-depressed adolescent volunteers, collected using a human-computer interaction system.
-Self-reporting anxiety and depression rating scales were also collected and verified using the same system, which can be used as gold standard in training and evaluating machine-learning based techniques.
-After rigorous data preprocessing and cleaning, the corpus comprises 7,758 interview data points for anxiety detection and 4,266 for depression detection.
-To the best of our knowledge, MMPsy is the first publicly available adolescent psychological assessment corpus that simultaneously detects anxiety and depression while providing both audio and text data in Chinese.
-In addition, we propose a novel mental disorder estimation network, termed Mental-Perceiver, to automatically detect anxious or depressive mental states based on users’ audio inputs and corresponding transcripts.
-The Mental-Perceiver leverages attention mechanisms to map multimodal inputs and category semantic priors to a fixed-size multimodal feature, using a learnable, sharable embedding for effective multimodal fusion.
-It then applies a deep, fully attentional network to process the fused multimodal feature for in-depth multimodal feature extraction.
-Finally, the network decodes the extracted feature using a learnable query array to produce the final mental state estimation.
-Extensive experiments conducted on the MMPsy corpus and the public DAIC-WOZ dataset demonstrate the effectiveness and superiority of our proposed Mental-Perceiver model.
-Our contributions can be summarized as follows:
-  • We introduce MMPsy, a large-scale multimodal psychological assessment corpus containing 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection.
-  • We develop Mental-Perceiver, a novel mental disorder estimation network that automatically detects anxious or depressive mental states from audio and transcript data, utilizing a fully attentional network architecture for effective multimodal fusion.
-  • Extensive experimental results on MMPsy and the public DAIC-WOZ dataset validate the effectiveness and superiority of our proposed Mental-Perceiver model in automatic mental disorder detection.
+2014), are restricted in scale and scope.
+Although initiatives like EATD-Corpus (Shen, Yang, and Lin 2022) have been undertaken using self-reported mental health questionnaires, challenges persist regarding data size and validation.
+In this work, our goal is to address the challenges in facilitating automatic detection of anxiety and depression and to extend the research boundaries to a more real world setting.
+First, we introduce a large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in Mandarin-speaking adolescents.
+The MMPsy comprises audio and extracted transcripts of responses from anxious/depressed and non-anxious/non-depressed adolescent volunteers.
+After data pre-processing and cleaning, 7,736 cleaned interview data are available for anxiety detection and 4,247 for depression detection.
+To our knowledge, MMPsy is the first publicly available adolescent psychological assessment corpus capable of simultaneously detecting anxiety and depression, and containing both audio and text data in Chinese.
+Subsequently, we propose a deep learning model, termed Mental-Perceiver, to automatically detect anxious/depressive mental states based on users’ audio and corresponding transcripts.
+Mental-Perceiver first maps multi-modal inputs and category semantic priors to a fixed-size feature space with learnable embeddings for multi-modal fusion.
+It then employs a fully attentional network (Jaegle et al.
+2021) to further process the fused multimodal features and decode the underlying mental state with a learnable query array.
+Extensive experiments on our MMPsy and DAIC-WOZ datasets demonstrate the effectiveness and superiority of our proposed Mental-Perceiver model for the estimation of anxiety and depression.
+The key contributions of this work are summarized as follows:
+  • We introduce MMPsy, a large-scale multi-modal corpus for mental health assessment.
+  • We propose Mental-Perceiver, a fully attentional network for automated anxiety/depression detection from audio and transcripts.
+  • Extensive experiments on the MMPsy and DAIC-WOZ datasets demonstrate the effectiveness of the Mental-Perceiver.
 Related Work
+Anxiety/Depression Detection Datasets
+Publicly available datasets for the detection of anxiety and depression remain limited due to the challenges of collecting sensitive mental health data.
+Existing datasets, such as DAIC-WOZ (Gratch et al.
+2014) and AViD-Corpus (Valstar et al.
+2014), focus primarily on depression and often lack textual transcripts or sufficient sample sizes.
+The MODMA dataset (Cai et al.
+2022), while incorporating EEG signals along with audio and video recordings, presents challenges for real-world application due to the difficulty in collecting EEG data in regular scenarios and its limited sample size.
+Recent efforts like EATD-Corpus (Shen, Yang, and Lin 2022) have attempted to address these limitations by incorporating self-reported measures and textual data, but are constrained by a relatively small sample size.
+This scarcity of diverse and large-scale datasets hinders the development and evaluation of robust machine learning models for anxiety and depression detection, particularly in multilingual and cross-cultural contexts.
 Automatic Anxiety/Depression Detection
-Early studies of automatic anxiety/depression detection were concentrated on extracting effective features from the responses to questions that were highly correlated with anxiety/depression.
+Early research in automatic anxiety/depression detection focused on extracting relevant features from interview responses.
 Williamson et al.
 (Williamson et al.
-2016) used related semantic context cues entailed in the voice, video-based facial action units, and transcribed text of individuals and built a Gaussian Staircase Model to detect depression automatically.
+2016) utilized semantic cues from voice, facial action units, and transcribed text within a Gaussian Staircase Model for depression detection.
 Yang et al.
 (Yang et al.
-2016) selected depression-related questions after analyzing interview transcripts manually and constructed a decision tree with the selected questions to predict the participants’ depression states.
+2016) employed manual transcript analysis to select depression-related questions and built a decision tree for prediction.
 Similarly, Sun et al.
 (Sun et al.
-2017) first manually selected questions related to certain topics such as recent feelings and sleep quality by conducting content analysis based on the text transcripts of clinical interviews.
-Then, they extracted text features from these selected questions as the input of Random Forest to detect depression tendencies.
+2017) used content analysis to select text features from interview transcripts and applied Random Forest to detect depression tendency.
 Gong et al.
-(Gong and Poellabauer 2017) performed topic modeling to split the interviews into topic-related segments.
-Then, they maintained the most discriminating features by a feature selection algorithm.
-Giannakakis et al.
+(Gong and Poellabauer 2017) utilized topic modeling for feature selection, while Giannakakis et al.
 (Giannakakis et al.
-2017) focused on non-voluntary and semi-voluntary facial cues to estimate the emotion representation more objectively.
-They selected the most robust features from the features including eye-related events, mouth activity, head motion parameters, and heart rate estimated through camera-based photoplethysmography.
-Finally, they deployed a ranking transformation to investigate the correlation of facial parameters with a participant’s perceived amount of stress/anxiety.
-With the advancement of deep learning, extracting and integrating multi-modal features through deep learning models have become particularly promising for anxiety/depression detection.
+2017) investigated the correlations between facial cues and perceived stress/anxiety.
+Deep learning advancements have enabled more effective multi-modal feature extraction and integration.
 Ma et al.
 (Ma et al.
-2016) encoded the depressive audio characteristics with the combination of CNN and LSTM to predict the presence of depression.
+2016) used CNN-LSTM models for depressive audio encoding.
 Yang et al.
 (Yang et al.
-2017) trained a deep Convolution Neural Network (CNN)-based depression detection model with specially-designed a set of audio and video descriptors.
+2017) used deep CNNs with audiovisual descriptors for the detection of depression.
 Tuka et al.
-(Al Hanai, Ghassemi, and Glass 2018) selected audio features and text features that were strongly related to depression severity by computing Pearson Coefficients and built a Long Short-Term Memory (LSTM) network to assess depression tendency.
+(Al Hanai, Ghassemi, and Glass 2018) leveraged Pearson coefficients to select relevant audio and text characteristics for the evaluation of depression based on LSTM.
 Haque et al.
 (Haque et al.
-2018) proposed a causal CNN model to summarize acoustic, visual, and linguistic features into embeddings which were then used to predict depressive states.
+2018) proposed a causal CNN to generate embeddings from acoustic, visual, and linguistic features for the prediction of depression.
 Shen et al.
-(Shen, Yang, and Lin 2022) proposed a depression detection approach utilizing speech characteristics and linguistic contents from participants’ interviews.
-Lin et al.
+(Shen, Yang, and Lin 2022) and Lin et al.
 (Lin et al.
-2022) used the biological information of speech, combined with deep learning to build a rapid binary classification model of depression in the elderly.
-Agarwal (Agarwal, Jindal, and Singh 2023) developed machine-learning solutions to diagnose anxiety disorders from audio journals of patients.
-Anxiety/Depression Detection Datasets
-The public anxiety/depression datasets are quite scarce (Gratch et al.
-2014; Valstar et al.
-2014; Shen, Yang, and Lin 2022) due to ethical issues.
-To the best of our knowledge, there are only three publicly available datasets  (Gratch et al.
-2014; Valstar et al.
-2014; Shen, Yang, and Lin 2022) referring to depression detection while there are no publicly available multi-modal datasets including audio, facial, or text referring to anxiety detection.
-The DAIC-WoZ dataset (Gratch et al.
-2014) contains recordings and transcripts of 142 American participants who were clinically interviewed by a computer agent automatically.
-AVid-Corpus (Valstar et al.
-2014) contains audios and videos of German participants answering a set of queries or reciting fables.
-However, the text transcripts are not provided.
-EATD-Corpus (Shen, Yang, and Lin 2022) is released to facilitate the research in depression detection.
-It consists of audios and text transcripts extracted from the interviews of 162 student volunteers recruited from Tongji University.
-Although EATD-Corpus provides the text transcripts, its data scale is small.
+2022) focused on speech and linguistic content for depression detection, while Agarwal (Agarwal, Jindal, and Singh 2023) explored machine learning for anxiety diagnosis from audio journals.
+Recent years have witnessed significant advances in transformer-based and LLM approaches for the detection of mental disorders.
+Ji et al.
+(Ji et al.
+2021) introduced domain-specific models, MentalBERT and MentalRoBERTa, which demonstrated superior performance on mental health benchmarks compared to general-purpose LLMs.
+Xu et al.
+(Xu et al.
+2023) revealed that instruction-tuned smaller LLMs could achieve higher balanced accuracy than larger models in mental health prediction tasks, while Binz and Schulz (Binz and Schulz 2023) enhanced LLMs’ emotional state modeling capabilities by fine-tuning with psychological experiment data.
+In parallel, Qi et al.
+(Qi et al.
+2023) achieved notable improvements in cognitive distortion and suicide risk detection using fine-tuned GPT-3.5 models.
+The field has increasingly embraced multi-modal approaches, as evidenced by Ahmed et al.
+(Ahmed et al.
+2023), who developed a comprehensive framework that incorporates video, audio, text and EEG data with uncertainty management capabilities.
+Building on this trend, Ding et al.
+(Ding et al.
+2024) proposed IntervoxNet, a dual-modal audio-text fusion model using transformer architectures, achieving an impressive F1 score of 0.90 in the detection of depression from interview data.
 MMPsy: A New Benchmark
-Since the public anxiety/depression datasets are quite scarce, we construct and release a new Chinese multi-modal mental health dataset named MMPsy to facilitate the research in both anxiety detection and depression detection.
-Like EATD-Corpus (Shen, Yang, and Lin 2022), MMPsy consists of audios and text transcripts extracted from the interviews of more than 20 thousand primary and secondary school students in the Guangdong Province of China.
-All the volunteers have signed informed consent and tried their best to guarantee the authenticity of all the information provided.
-Each volunteer is required to answer 10 specially designed questions about anxiety and depression and complete a GAD-7 questionnaire (Mossman et al.
-2017) or PHQ-9 questionnaire (Kroenke, Spitzer, and Williams 2001) for different disorders detection.
-The GAD-7 questionnaire consists of 7 items that measure the severity of anxiety.
-It is a commonly used questionnaire for psychologists to screen the anxiety severity of individuals in practice.
-A raw GAD-7 score can be summarized from the questionnaire.
-For Chinese people, an index GAD-7 score greater than or equal to 10 implies that the individual is in anxiety.
-Similarly, the PHQ-9 questionnaire consists of 9 items that measure the severity of depression.
-It is a commonly used questionnaire for psychologists to screen the depression severity of individuals.
-A raw PHQ-9 score can be summarized from the questionnaire.
-For Chinese people, an index PHQ-9 score greater than or equal to 10 implies that the individual is in depression.
-According to the criterion, there are 704 anxious volunteers and 7,032 non-anxious volunteers in the anxiety subset of MMPsy while there are 853 depressed volunteers and 3,394 non-depressed volunteers in the depression subset of MMPsy.
-In the final dataset, the overall duration of response audios about anxiety is about 145.9 hours and the overall duration of response audios about depression is about 69.5 hours.
+Characteristic Anxiety Subset Depression Subset
+Gender
+Males 2,262 (53.02%) 4,087 (52.68%)
+Females 2,004 (46.98%) 3,671 (47.32%)
+Grade Level
+Grade 4 335 -
+Grade 5 987 -
+Grade 6 1,342 -
+Grade 7 1,646 1,564
+Grade 8 1,554 1,397
+Grade 9 1,579 1,305
+Grades 10-12 315 -
+Table 1: Demographic Information of Data Collection Participants
+Data collection was conducted during the mandated annual mental health assessment of elementary and middle school students in Guangdong Province, China, as required by local government regulations [^1] Guangzhou Municipal Education Bureau Website - Regulations on Promoting Mental Health of Primary and Secondary School Students in Guangzhou (Full Text): https://jyj.gz.gov.cn/yw/jyyw/content/post˙9907147.html.
+The research team received approval from schools and their administrative Education Bureaus to provide mental health assessment services and collect data for research purposes.
+Under institutional supervision, data were collected from over 10,000 primary and secondary school students.
+All participants provided informed consent with guidance from supervising teachers or social workers.
+Demographic characteristics of the participants are presented in Table 1.
+Participants were asked to provide spontaneous responses to 10 specifically designed questions addressing anxiety and depression.
+Additionally, they completed either the GAD-7 (Mossman et al.
+2017) or PHQ-9 (Kroenke, Spitzer, and Williams 2001) questionnaire for respective disorder detection.
+The GAD-7, comprising seven items measuring anxiety severity, is a standardized screening tool widely used in clinical practice.
+The PHQ-9, consisting of nine items measuring depression severity, serves a similar function for depression screening.
+For Chinese populations, scores $\geq$ 10 on either scale indicate the presence of the respective disorder.
+Based on these criteria, the anxiety subset of MMPsy includes 704 anxious and 7,032 non-anxious participants, while the depression subset comprises 853 depressed and 3,394 non-depressed participants.
 Parts Subset $\times$ $\checkmark$ Avg Duration (Sec)
@@
 Table 2: Statistics for our MMPsy dataset. $\times$ means the label is non-anxious or non-depressed and $\checkmark$ denotes the data label is anxious or depressed.
-The construction of MMPsy can be summarized as two steps: data collection and data preprocessing.
+The construction of MMPsy comprised two primary phases: data collection and preprocessing.
   • Data collection.
-We develop a web app to conduct the interview and to collect audio responses and the questionnaires.
-Our web app will ask the interviewees 10 questions and let them finish the GAD-7 questionnaires or PHQ-9 questionnaires.
-During the interview, the audio responses will be recorded and uploaded to the server automatically.
-The results of the questionnaires also will be uploaded.
-In this step, we collected 17,247 raw interviews about anxiety and 11,306 raw interviews about depression.
-To ensure the authenticity of the questionnaires, we added some extra but irrelevant questions in the questionnaires, which are used to check the response reliability.
+A custom web application was developed to conduct interviews and collect audio responses and questionnaire data.
+The application administered 10 questions along with the GAD-7 or PHQ-9 questionnaires.
+The audio responses were automatically recorded and uploaded to the server along with the questionnaire results.
+This phase yielded 17,247 raw interviews for anxiety and 11,306 for depression.
+To ensure the validity of the responses, additional control questions were embedded within the questionnaires to assess the reliability of the responses.
   • Data preprocessing.
-Since the raw interviews are noised, we perform several preprocessing operations on the collected data.
-First, we filtered out the unauthentic data by checking the responses of those extra questions in the questionnaires.
-Second, mute audios and audios less than 1 second are removed.
-Besides, we also removed the silent segments identified with voice activity detection in each recording.
-Third, the background noises are eliminated using Spleeter (Hennequin et al.
+The raw interview data underwent several preprocessing steps.
+Initially, responses were filtered based on the control questions to eliminate inauthentic entries.
+Subsequently, mute recordings and those shorter than 1 second were removed, and silent segments were eliminated using voice activity detection.
+Background noise was then removed using Spleeter (Hennequin et al.
 2020).
-After that, Paraformer (Gao et al.
-2022) is deployed to extract textual transcripts from audios.
-To ensure the correctness of extracted transcripts, we checked and corrected the transcript manually by listening to the audio and comparing the consistency between the semantics of auidos and the transcripts at the word level.
-After data preprocessing, we obtained 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection.
-We randomly split these data into a training set, a validation set, and a test set in an 8:1:1 ratio.
-In this way, the anxiety detection part of MMPsy consists of 6,188 training volunteers, 744 validation volunteers, and 744 test volunteers while the depression detection part of MMPsy contains 3,397 training volunteers, 425 validation volunteers, and 425 test volunteers.
-The data statistics are listed in the Table 1.
-For each volunteer, we pack their audios and transcripts according to their sequential order and resegment the data in 60 seconds with a 10-second overlap like the way in the work (Wei et al.
-2022).
-Thus, the data of a volunteer can generate multiple samples.
+Textual transcripts were extracted using Paraformer (Gao et al.
+2022) and manually verified through audio comparison to ensure semantic consistency at the word level.
+  • Data anonymization and reporting.
+To protect participant privacy, all sensitive information and any data that could potentially identify participants were meticulously removed during the preprocessing stage.
+Anonymization was carried out in compliance with ethical guidelines and data protection regulations.
+Additionally, for students whose responses on the GAD-7 or PHQ-9 scales indicated potential anxiety or depression problems, their data was securely reported to the designated school administrative offices, as required by institutional and regulatory policies, ensuring appropriate follow-up support.
+The preprocessing phase yielded 7,736 cleaned interviews for anxiety detection and 4,247 for depression detection.
+These data were randomly partitioned into training, validation, and test sets using an 8:1:1 ratio.
+Consequently, the anxiety detection subset comprises 6,188 training, 774 validation, and 774 test participants, while the depression detection subset contains 3,397 training, 425 validation, and 425 test participants.
+The complete data statistics are presented in Table 2.
+Following the methodology of (Wei et al.
+2022), each participant’s audio and transcript data were sequentially organized and segmented into 60-second intervals with 10-second overlaps, generating multiple samples per participant.
+Figure 1: Illustration of Mental-Perceiver.
 Mental-Perceiver
-Figure 1: Illustration of Mental-Perceiver.
 This section describes the architecture of our proposed Mental-Perceiver as shown in Figure 1.
-We first encode by applying an attention module that maps multimodal input $x\in\mathbb{R}^{M\times D_{x}}$ to features in a latent space $z\in\mathbb{R}^{2\times D_{z}}$ by interacting with category prior $p\in\mathbb{R}^{2\times D_{p}}$ which is obtained by computing the center point of different text representations from different categories.
-Then, we conduct deep feature extraction on latent features $z$ by applying a series of attention modules that take in and return features $z^{\prime}$ in this latent feature space.
+We first encode by applying an attention module that maps multimodal input $x\in\mathbb{R}^{M\times D_{x}}$ to features in a latent space $z\in\mathbb{R}^{2\times D_{z}}$ by interacting with a category prior $p\in\mathbb{R}^{2\times D_{p}}$ which is obtained by computing the center point of different text representations from different categories.
+Then, we perform deep feature extraction on latent features $z$ by applying a series of attention modules that take in and return features $z^{\prime}$ in this latent feature space.
 Finally, we decode by applying an attention module that maps latent arrays $z^{\prime}$ and the query array $q\in\mathbb{R}^{2\times D_{q}}$to the final feature representation $y\in\mathbb{R}^{2\times D_{y}}$.
-Based on the final feature representation $y$, we apply a Linear layer to map the feature $y$ into class-wise logit outputs $y^{C_{0}}\in\mathbb{R}^{1\times 2}$ and $y^{C_{1}}\in\mathbb{R}^{1\times 2}$.
-$M$ is the length of multimodal input while $D_{x}$, $D_{z}$, $D_{q}$, and $D_{y}$ denote the feature dimension.
+Based on the final feature representation $y$, we apply a linear layer to map the feature $y$ to class-wise logit outputs $y^{C_{0}}\in\mathbb{R}^{1\times 2}$ and $y^{C_{1}}\in\mathbb{R}^{1\times 2}$.
+$M$ is the length of multimodal input, while $D_{x}$, $D_{z}$, $D_{q}$, and $D_{y}$ denote the feature dimension.
 With class-wise logits $y^{C_{0}}$ and $y^{C_{1}}$, to predict the class $c^{\prime}$ of the input $x$, we compute the mean values of corresponding vector elements in the $y^{C_{0}}$ and $y^{C_{1}}$ to obtain final logits $y^{\prime}$ followed by a Softmax function.
@@
 Each attention module applies a global query-key-value (QKV) attention operation followed by a multi-layer perceptron (MLP).
-The MLP is independently applied to each element of the index dimension.
+The MLP is applied independently to each element of the index dimension.
 Both the encoder and decoder take in two input feature arrays.
-The first array is used as input to the attention module’s key and value networks, and another array is used as input to the module’s query network for interacting and fusing with the first array.
-The attention module’s output has the same index dimension (the same number of elements) as the query input.
+The first array is used as input to the attention module’s key and value networks, and another array is used as input to the module’s query network to interact and fuse with the first array.
+The output of the attention module has the same index dimension (the same number of elements) as the query input.
 Therefore, the attention module can be modeled as follows:
@@
 To optimize our Mental-Perceiver, we deploy two losses.
-The first one is the matching loss $\mathcal{L}_{match}$ and another one is the classification loss $\mathcal{L}_{cls}$.
+The first one is the matching loss $\mathcal{L}_{match}$ and another is the classification loss $\mathcal{L}_{cls}$.
 Both these two losses are binary cross-entropy loss functions.
@@
   • SVM (Pedregosa et al.
-2011): a robust shallow model capable of performing binary classification efficiently by using a kernel trick, representing the data only through a set of pairwise similarity comparisons between the original data points using a kernel function, which transforms them into coordinates in the higher dimensional feature space.
+2011): a robust shallow model capable of performing binary classification efficiently by using a kernel trick, which transforms original datapoints into coordinates in the higher dimensional feature space.
   • RandomForest (Pedregosa et al.
@@
   • PerceiverIO (Jaegle et al. 2021): it is a general-purpose architecture that handles multimodal data with fully attention design and a flexible querying mechanism.
+  • AFABNet (Xu et al. 2024): it is an attention-based acoustic feature fusion network for depression detection by combining four different acoustic features.
+  • Qwen2-Audio-Instruct (Chu et al.
+2024): it is a Qwen large audio-language model (7B) that can accept various audio signal inputs and perform audio analysis or direct textual responses with regard to speech instructions.
 For the shallow models SVM, RandomForest, and XGBoost, we provide the following audio features as input: F0 statistics (mean), log-energy, zero-crossing-rate, loudness, pitch period entropy, jitters, shimmers, harmonics-to-noise ratio, detrended fluctuation analysis, linear spectral coefficients-0, linear spectral frequencies-0, formants (F1), and amplitude Shannon entropy.
@@
 PerceiverIO 0.83 0.74 0.63 0.84 0.95 0.84 0.90 0.29 0.63 0.40
+AFABNet 0.82 0.63 0.41 0.86 0.93 0.86 0.89 0.22 0.41 0.29
+Qwen2-Audio-Instruct 0.72 0.76 0.80 0.71 0.97 0.71 0.82 0.22 0.80 0.34
 Mental-Perceiver (Ours) 0.85 0.76 0.65 0.87 0.96 0.87 0.92 0.34 0.65 0.45
@@
 PerceiverIO 0.81 0.77 0.66 0.85 0.91 0.85 0.88 0.53 0.66 0.59
+AFABNet 0.78 0.58 0.24 0.92 0.83 0.92 0.87 0.44 0.24 0.31
+Qwen2-Audio-Instruct 0.79 0.58 0.21 0.94 0.82 0.94 0.88 0.45 0.21 0.29
 Mental-Perceiver (Ours) 0.85 0.79 0.69 0.89 0.92 0.89 0.9 0.61 0.69 0.64
@@
 PerceiverIO 0.7 0.58 0.29 0.88 0.74 0.88 0.81 0.5 0.29 0.36
+AFABNet 0.74 0.63 0.36 0.91 0.77 0.91 0.83 0.62 0.36 0.45
+Qwen2-Audio-Instruct 0.70 0.65 0.52 0.78 0.79 0.78 0.79 0.50 0.52 0.51
 Mental-Perceiver (Ours) 0.79 0.66 0.36 0.97 0.78 0.97 0.86 0.83 0.36 0.5
@@
 On the contrary, FP represents cases where the model incorrectly labeled instances as positive, and FN denotes instances of the positive class that were incorrectly classified as negative.
-Based on these concepts, several common performance metrics can be defined to assess the model’s performance:
+Based on these concepts, several common performance metrics are adopted to assess the model’s performance:
   • Accuracy: Accuracy (Acc) represents the proportion of all predictions that are correctly predicted.
@@
   • Specificity: Specificity (Spec) is the ratio of negative predictions to the number of actual negatives, and therefore identical to Recall(0).
-  • UAR: Although Acc is often used to evaluate model performance, it suffers from the data imbalance issue.
-The stronger the imbalance the more accuracy tends to reflect the performance of the majority class since it is a weighted accuracy.
-For this reason in areas where relatively small data sets predominate, such as the bio-medical field or in the paralinguistics field in general, researchers usually prefer and report the Unweighted Average Recall (UAR), which is defined as the average of Sensitivity and Specificity.
-    $UAR=\frac{Sensitivity+Specificity}{2}$ (8)
+  • UAR: Unlike accuracy, which is biased by data imbalance, Unweighted Average Recall (UAR) is preferred in fields with imbalanced datasets, such as biomedical and paralinguistics.
+UAR, i.e., the average of Sensitivity and Specificity, provides a more balanced performance measure.
 During these metrics, we use Accuracy (Acc), Unweighted Average Recall (UAR), Sensitivity (Sens), and Specificity (Spec) as the main evaluation metrics for evaluating overall model performance.
@@
 Main Results
-The experiment results of our Mental-Perceiver and various baselines on three datasets MMPsy-Anxiety, MMPsy-Depression, and DAIC-WOZ are shown in Table 2.
-From the results on different datasets, we can conclude as follows.
-On the MMPsy-Anxiety dataset, our Mental-Perceiver outperforms baselines on Acc, UAR, and Sens while achieving competitive performance on Spec.
+The experiment results of our Mental-Perceiver and various baselines on MMPsy-Anxiety, MMPsy-Depression, and DAIC-WOZ are shown in Table 3.
+From the results on different datasets, we can draw the following conclusions.
+1) On MMPsy-Anxiety, our Mental-Perceiver outperforms baselines on Acc and UAR while achieving competitive performance on Sens and Spec.
 This shows that our Mental-Perceiver can achieve better overall anxiety detection performance with a better trade-off between Sensitivity and Specificity.
 For different categories, we can observe that our Mental-Perceiver achieves relatively high performance in the normal category and achieves the best performance in the anxious category.
-On MMPsy-Depression dataset, our Mental-Perceiver outperforms baselines on Acc, UAR, and Sens while achieving competitive performance on Spec.
+2) On MMPsy-Depression, our Mental-Perceiver outperforms baselines on Acc, UAR, and Sens while achieving competitive performance on Spec.
 This shows that our Mental-Perceiver can achieve better overall depression detection performance with a better trade-off between Sensitivity and Specificity.
 For different categories, we can observe that our Mental-Perceiver achieves the best precision and F1-score in the normal category while outperforming all baselines in the depressed category on all three metrics.
-On DAIC-WOZ dataset, a similar conclusion can be reached.
+3) On DAIC-WOZ, a similar conclusion can be reached.
 The Mental-Perceiver outperforms baselines on Acc, UAR, and Spec.
@@
 Conclusion
-In this work, we construct a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in adolescents who use Mandarin.
+In this work, we construct a new large-scale Multi-Modal Psychological assessment corpus (MMPsy) about anxiety and depression in adolescents who speak Mandarin.
 The MMPsy contains audios and extracted transcripts of responses from anxious/depressed and non-anxious/non-depressed adolescent volunteers.
-There are 7,758 cleaned interview data for anxiety detection and 4,266 for depression detection.
+There are 7,736 cleaned interview data for anxiety detection and 4,247 for depression detection.
 We further propose a novel mental disorder estimation network, named Mental-Perceiver, to detect anxious/depressive mental states automatically according to users’ audio and corresponding transcripts.
 Extensive experiments on MMPsy and the public DAIC-WOZ show the effectiveness and superiority of our proposed Mental-Perceiver.
+Acknowledgements
+This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No.
+62206314, GuangDong Basic and Applied Basic Research Foundation under Grant No.
+2022A1515011835, Science and Technology Projects in Guangzhou under Grant No.
+2024A04J4388 and China Postdoctoral Science Foundation funded project under Grant No.
+2021M703687.